Will you please explain a step by step of how to answer this question:
Kristin is driving on the highway at 60 miles per hour. How long will it take Kristin to reach a rest area 90 miles away?

Answer by Alan3354(69443)

Kristin is driving on the highway at 60 miles per hour. How long will it take Kristin to reach a rest area 90 miles away?
===============
distance = rate*time
or,
rate = d/t
To get miles per hour, you divide the miles by the hours.
--------------
distance = rate*time
d = r*t
Going 60 mi/hr, you go 60 miles in an hour. Is that not obvious?
Then, in another 1/2, you go another 30 miles. 1/2 the time, 1/2 the distance.
--> 90 miles in 1.5 hours
==================
t = d/r
t = 90 mi/60 mi/hr
t = 1.5 hours
=====================
That's a lot to say about such a simple concept.